OpenStudy (anonymous):

calculate y' if y= 10^tanpitheta please help and show all steps!

OpenStudy (anonymous):

\[y=10^{\tan(\pi x)}\] \[y'=\ln(10)\times 10^{\tan(\pi x)}\frac{d}{dx}\tan(\pi x)\] is s start

OpenStudy (anonymous):

\[y'=10^{\tan (\pi \theta)}*\ln10*\pi \sec ^{2}(\pi \theta)\]

OpenStudy (anonymous):

\[y=a ^{f(x)} \] a= reel number then \[y'=a ^{f(x)}*lna*f(x)'\]